​Q. A wire has radius of 0.50mm and carries a current 0.5A. The Resistivity of the material of wire is 1.1*10^-6 ohm-m. The potential difference per unit length along the wire is ???????

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
The potential difference is I*R, where R is the wire's resistance. R = ρ*L/A where
ρ = resistivity
L = length
A = area
ρ = 1.1*10^-6 ohm-m
L = 1 m (unit length)
A = π*d²/4 = π*(5*10^-4)²/4 = 1.96*10^-7 m²
R = 5.61 ohm/m
E = 0.50*5.61 = 2.805 v/m
 
Regards
Arun (askIITians forum expert)